OpenStudy (anonymous):

∑[k=1,4](-(1)/(3))^k find the indicted sum

OpenStudy (anonymous):

(-(1)/(3))^1 + (-(1)/(3))^2 + (-(1)/(3))^3 = (-(1)/(3))^4 = -1/3 + 1/9 - 1/27 + 1/81 = -20/81 according to my calculator

OpenStudy (anonymous):

Oops change that = in my first line to +